calculate the mass ( in kg ) of chlorine in 70 kg CFCl3
Answer #1

Molar mass of CFCl3,

MM = 1*MM(C) + 1*MM(F) + 3*MM(Cl)

= 1*12.01 + 1*19.0 + 3*35.45

= 137.36 g/mol

mass(CFCl3)= 70 Kg

= 70000 g

use:

number of mol of CFCl3,

n = mass of CFCl3/molar mass of CFCl3

=(7*10^4 g)/(1.374*10^2 g/mol)

= 5.096*10^2 mol

This is number of moles of CFCl3

one mole of CFCl3 contains 3 mole of Cl

use:

number of moles of Cl = 3 * number of moles of CFCl3

= 3 * 5.096*10^2

= 1.529*10^3

Molar mass of Cl = 35.45 g/mol

use:

mass of Cl,

m = number of mol * molar mass

= 1.529*10^3 mol * 35.45 g/mol

= 5.42*10^4 g

= 54.2 Kg

Answer: 54.2 Kg